# arXiv PreprintSubmitted to arXiv on August 8, 2025

### v3

@@ Methods.
Figure LABEL:fig:timing-curves shows separate training and prediction times. Training with scikit-learn scales cubically in $m$; prediction scales quadratically, but still rises steeply for large $m$. In contrast, EnKF and LETKF training scale linearly in ensemble size and prediction is effectively $O(1)$ for fixed ensemble, yielding two nearly flat curves. As the problem size scales beyond feasibility of GP regression, we expect to achieve efficiency gains by adopting mature DA techniques.
[…]
fig:posterior-samples Refer to caption
Figure 1: Posterior draws from the exact GP (sklearn), EnKF, and LETKF..
@@
The results confirm that a standard EnKF is a computationally efficient method for path-wise GP sampling. More importantly, the equivalence acts as a bridge to a rich ecosystem of data assimilation techniques. The LETKF demonstrates this powerfully: by simply changing one line of code to invoke a localized filter, we gain access to a highly scalable inference method that parallels sparse or domain-decomposition approaches in the GP literature. This opens the door for the GP community to leverage decades of mature, practical DA engineering “for free.”
@@ 9 Conclusion and Implications
The literature of the EnKF is vast, and we have only scratched the surface of the many variants and extensions that have been proposed. Many of the ideas developed for the EnKF are thus potentially applicable to Gaussian Process regression via the Matheron update. For example, it provides a clear probabilistic interpretation of the ensemble update that may inspire improved regularization and localization strategies. For one example, the Local Ensemble Transform Kalman Filter (Bocquet et al., 2020) provides state-of-the-art performance in high-dimensional filtering and system identification settings, and could be adapted to the Matheron update.
@@
The original of this contentwas drafted by human hands, with the exception of the experiments section and supporting code which was generated by LLMs.

### v4

# arXiv PreprintSubmitted to arXiv on September 18, 2025

@@ Methods.
Figure LABEL:fig:timing-curves reports fit (training) and predict times as we sweep (top) the number of observations $m$ at fixed state dimension $d$, and (bottom) the state dimension $d$ at fixed $m$. The exact GP (scikit-learn) shows the expected $\mathcal{O}(m^{3})$ fit scaling. Its predict time grows with $d$ and mm—mean-only scales as $\mathcal{O}(d\,m)$, while computing both mean and standard deviation (our default) scales as $\mathcal{O}(d\,m^{2})$. In contrast, EnKF and LETKF fits scale roughly linearly in $m$ for fixed ensemble size $N$ (and linearly in $d$ with localization), while prediction is linear in $d$ to read out the analysis mean (and $\mathcal{O}(Nd)$ if we also materialize an ensemble). We do not vary $N$ in these plots. Consequently, as $d$ grows large at fixed $m$, the DA methods yield flatter predict‑vs‑$d$ curves than the exact GP, and as $m$ grows at fixed $d$, DA fits avoid the cubic blow‑up of exact GP regression.
Figure 1: Posterior draws from the exact GP (sklearn), EnKF, and LETKF.
@@
The results confirm that a standard EnKF is a computationally equivalent method for path-wise GP sampling. More importantly, the equivalence acts as a bridge to a rich ecosystem of data assimilation techniques. The LETKF demonstrates this: by simply changing one line of code to invoke a localized filter, we gain access to a highly scalable inference method that parallels sparse or domain-decomposition approaches in the GP literature. This opens the door for the GP community to leverage decades of mature, practical DA engineering “for free.”
@@ 9 Conclusion and Implications
The literature of the EnKF is vast, and we have only scratched the surface of the many variants and extensions that have been proposed. Many of the ideas developed for the EnKF are thus potentially applicable to Gaussian Process regression via the Matheron update. For example, it provides a clear probabilistic interpretation of the ensemble update that may inspire improved regularization and localization strategies.
@@
This content was drafted by human hands, with the exception of the experiments and supporting code which was generated by LLMs.